OpenStudy (uri):

In how many ways can 5 people be seated on a sofa,If there are only 3 seats available.

OpenStudy (anonymous):

This is how you do it. Since there are 3 seats available and only 5 people you multiply like this: 5*4*3 = ?. You go down 1 number for each time you multiply, so there are no duplicates, and only multiply like that for 3 times, because that is the number of seats available.

OpenStudy (uri):

I got half of what you said.

OpenStudy (nincompoop):

Combinations without repetition (n=5, r=3) Using the first 5 items: {a,b,c,d,e} List has 10 entries. {a,b,c} {a,b,d} {a,b,e} {a,c,d} {a,c,e} {a,d,e} {b,c,d} {b,c,e} {b,d,e} {c,d,e}

OpenStudy (uri):

How @nincompoop

OpenStudy (nurali):

5P3 = 5!/(5-3)! = 5*4*3 = 60 ways

OpenStudy (uri):

60 is the answer right?

OpenStudy (anonymous):

Yes

OpenStudy (selrachcw95):

Uri this should go to math.

OpenStudy (selrachcw95):

Not history.

OpenStudy (nincompoop):

\[\frac{ n! }{ \left( n-r \right)!\left( r \right)! }\]

OpenStudy (uri):

I got it,I did that Ques ^10P3 it is done in same way,I got it..Wait nin you're doing in the method of 5C3?

OpenStudy (uri):

5C3 =5P3/3!

OpenStudy (uri):

Omg i posted in History,Dahell i didn't even see it :o @selrachcw95

OpenStudy (nincompoop):

that is without repetition

OpenStudy (uri):

So we have to do it in which method..? The one @mangorox and @Nurali said right? :)

OpenStudy (selrachcw95):

Np just close it and move to math. :D

OpenStudy (uri):

now i'm between and it's almost done. @selrachcw95

OpenStudy (uri):

I'm in*

OpenStudy (nincompoop):

if you want the order to matter then 60 using \[\frac{ n! }{ (n-r)! }\]

OpenStudy (selrachcw95):

ok since I'm extremely nice won't report u. :D

OpenStudy (nincompoop):

Permutations without repetition (n=5, r=3) Using the first 5 items: {a,b,c,d,e} List has 60 entries. {a,b,c} {a,b,d} {a,b,e} {a,c,b} {a,c,d} {a,c,e} {a,d,b} {a,d,c} {a,d,e} {a,e,b} {a,e,c} {a,e,d} {b,a,c} {b,a,d} {b,a,e} {b,c,a} {b,c,d} {b,c,e} {b,d,a} {b,d,c} {b,d,e} {b,e,a} {b,e,c} {b,e,d} {c,a,b} {c,a,d} {c,a,e} {c,b,a} {c,b,d} {c,b,e} {c,d,a} {c,d,b} {c,d,e} {c,e,a} {c,e,b} {c,e,d} {d,a,b} {d,a,c} {d,a,e} {d,b,a} {d,b,c} {d,b,e} {d,c,a} {d,c,b} {d,c,e} {d,e,a} {d,e,b} {d,e,c} {e,a,b} {e,a,c} {e,a,d} {e,b,a} {e,b,c} {e,b,d} {e,c,a} {e,c,b} {e,c,d} {e,d,a} {e,d,b} {e,d,c}

OpenStudy (uri):

Okay we can do it in both ways? @nincompoop :D

OpenStudy (selrachcw95):

Just remember where u are next time. :D

OpenStudy (uri):

@selrachcw95 Thanks :D but i didn't do it puposely :3

OpenStudy (selrachcw95):

Ok. :D

OpenStudy (selrachcw95):

I barley report anyone.Lol!!!

OpenStudy (nincompoop):

sometimes there are some rules or constraints that may be imposed like can we repeat numbers or should the order of number matter the first one I gave you cannot have the elements in the set repeated and the order didn't matter. so the formula is different.

OpenStudy (nincompoop):

@selrachcw95 just stop typing. you aren't even helping are you drunk or something?

OpenStudy (nincompoop):

oldrin is typing, now I am scared LUL

OpenStudy (nincompoop):

if you want numbers, this might be better Permutations without repetition (n=5, r=3) List has 60 entries. {1,2,3} {1,2,4} {1,2,5} {1,3,2} {1,3,4} {1,3,5} {1,4,2} {1,4,3} {1,4,5} {1,5,2} {1,5,3} {1,5,4} {2,1,3} {2,1,4} {2,1,5} {2,3,1} {2,3,4} {2,3,5} {2,4,1} {2,4,3} {2,4,5} {2,5,1} {2,5,3} {2,5,4} {3,1,2} {3,1,4} {3,1,5} {3,2,1} {3,2,4} {3,2,5} {3,4,1} {3,4,2} {3,4,5} {3,5,1} {3,5,2} {3,5,4} {4,1,2} {4,1,3} {4,1,5} {4,2,1} {4,2,3} {4,2,5} {4,3,1} {4,3,2} {4,3,5} {4,5,1} {4,5,2} {4,5,3} {5,1,2} {5,1,3} {5,1,4} {5,2,1} {5,2,3} {5,2,4} {5,3,1} {5,3,2} {5,3,4} {5,4,1} {5,4,2} {5,4,3}

OpenStudy (anonymous):

In how many ways can 5 people be seated on a sofa,If there are only 3 seats available. You want the number of ways that you can *permute* the people among 3 seats... consider each seat separately: